# Qualification-free convex analysis via the joint supporting subspace

### v4

@@ 3 Preliminaries
Intuitively the notion of a face of a convex set $C$ is a subset of $C$ for which there is no line segment in $C$ that “goes through” $F$. The following definitions can be found inline in [schneider_ConvexBodiesBrunn_2013, p. 16] and [schneider_ConvexBodiesBrunn_2013, p.75] respectively.
Definition 3.1 (Face of a convex set).
Let $C$ be convex. A subset $F\subseteq C$ is defined to be a face of $C$ when
$\forall y,z\in C,(\exists x\in]y,z[\cap F\implies y,z\in F).$
Definition 3.2 (Exposed face of a convex set).
Let $C\subseteq\mathbb{R}^{n}$ be convex. A subset $F\subseteq C$ is defined to be an exposed face of $C$ when
$\exists u\in\mathbb{R}^{n}\setminus\{0\}:F=\operatorname*{\arg\max}_{c\in C}\langle c,u\rangle.$
@@
For $x\in\mathbb{R}^{n}$ and a convex set $C\subseteq\mathbb{R}^{n}$, the normal cone of $C$ at $x$ is
$N_{C}(x)=\{y\mid\forall c\in C,\langle y,c-x\rangle\leq 0\}$
if $x\in C$, and $N_{C}(x)=\emptyset$ otherwise. We also recall some objects from convex analysis. Let $\bar{\mathbb{R}}=\mathbb{R}\cup\{-\infty,\infty\}$. For a convex function $f:\mathbb{R}^{n}\to\bar{\mathbb{R}}$, the subdifferential is
@@
with $\partial f(x)=\emptyset$ when $x\notin\operatorname{dom}(f)$. The convex conjugate is
$f^{*}(y):=\sup_{x\in\mathbb{R}^{n}}\langle y,x\rangle-f(x).$
@@ 4 The joint supporting subspace
The Minkowski sum $(+C\cap D)$ in the definition of $T_{a}$ is in fact an affine translation. This follows from the next corollary.
@@
We provide two other characterizations that follow from Theorem 4.1.
@@
It was shown in [rockafellarConvexAnalysis1970, Theorem 11.3] that this notion is equivalent to a lack of proper separation of the domains. We define the notions of separation, and proper separation below. We define these notions in affine subspaces, an intuitive generalization of separation in Euclidean spaces.
@@
Proper separation excludes the case where the two sets are merely separated by virtue of being both contained in a single hyperplane. Since separation follows from proper separation, and lack of proper separation is equivalent to intersection of relative interiors (by [rockafellarConvexAnalysis1970, Theorem 11.3]), the contrapositive statement is that
@@
The result follows from Theorem 4.2 by Equation 5, with the caveat that we defined separation in affine subspaces, which are not necessarily vector spaces, while in [rockafellarConvexAnalysis1970, theorem 11.3] this notion is defined in $\mathbb{R}^{n}$. Yet the argument still holds because the notions of separation, proper separation, and relative interiors are invariant under translation of $T_{a}$ to $T$. For any point $x_{0}\in T_{a}$, $C\cap T_{a}$ is (properly) separated from $C\cap T_{a}$ iff $C\cap T_{a}-x_{0}$ is (properly) separated from $D\cap T_{a}-x_{0}$ in $T$. A similar equivalence holds for intersection of relative interiors. ∎
@@ 5 Main results
With Corollary 4.5 we extend standard results of convex analysis by first localizing to $T_{a}$. We thus generalize the following beyond qualification conditions: subdifferential additivity; normal cones of the intersection of convex sets; attained infimal convolution for the conjugate of a sum $(f+g)^{*}$; the subdifferential chain rule; Fenchel-Rockafellar duality; and KKT conditions.
The next result implies that when standard qualification conditions hold, the classical formulations of our generalizations are trivially recovered.
@@ 5.1 Subdifferential calculus
By Corollary 4.5 applied to $\operatorname{dom}(f)$ and $\operatorname{dom}(g)$, the result follows by application of [rockafellarConvexAnalysis1970, Theorem 23.8] to the functions $f+\iota_{T_{a}}$ and $g+\iota_{T_{a}}$. ∎
@@
With Corollary 4.5 we can use [rockafellarConvexAnalysis1970, Corollary 23.8.1], from which we find that
@@
If $\operatorname{dom}(g)\cap\operatorname{\mathrm{range}}(A)=\emptyset,$ then the result holds trivially because $g(Ax)\equiv\infty$, and $(g+\iota_{T_{a}})(Ax)\equiv\infty$ as well. Otherwise, by Corollary 4.5, $\operatorname{relint}\operatorname{dom}(g+\iota_{T_{a}})\cap\operatorname{\mathrm{range}}(A)\neq\emptyset$, which lets us apply [rockafellarConvexAnalysis1970, Theorem 23.9] to the composition $g+\iota_{T_{a}}]\circ A$. Additionally, note that the only face of $\operatorname{\mathrm{range}}(A)$ is itself, whence by Lemma 4.1 $\operatorname{\mathrm{range}}(A)\subseteq T_{a}$. Since $0\in\operatorname{\mathrm{range}}(A)$, this in fact implies that $T_{a}=T$, and the result follows. ∎
@@ 5.2 Infimal convolution
Whenever $\operatorname{dom}(f)\cap\operatorname{dom}(g)=\emptyset$, we have that $T_{a}=\emptyset$, in which case the equality holds trivially, with the infimum attained at any $y\in\mathbb{R}^{n}$. Otherwise, by Corollary 4.5 the result follows from [rockafellarConvexAnalysis1970, Theorem 20.1]. ∎
@@ 5.3 Convex optimization
which is a weakening of Slater’s condition. A program with this property is called strongly consistent [rockafellarConvexAnalysis1970, Theorem 31.2].
@@
For $f:\mathbb{R}^{n}\to\bar{\mathbb{R}}$ and $g:\mathbb{R}^{m}\to\bar{\mathbb{R}}$ lower semicontinuous proper convex functions, a matrix $A\in\mathbb{R}^{n\times m}$, $T_{a}:=T_{a}(A\operatorname{dom}(f),\operatorname{dom}(g))$, and any $x\in\mathbb{R}^{n}$,
@@
where Equation 8 holds by [rockafellarConvexAnalysis1970, Theorem 6.6]; Equation 10 uses the identity that $A(S_{1}\cap A^{-1}(S_{2}))=AS_{1}\cap S_{2}$ for any sets $S_{1}\subseteq\mathbb{R}^{n}$, $S_{2}\subseteq\mathbb{R}^{m}$; and that the set in Equation 10 is non-empty is guaranteed by Corollary 4.5. ∎
@@
For $f:\mathbb{R}^{n}\to\bar{\mathbb{R}}$ and $g:\mathbb{R}^{m}\to\bar{\mathbb{R}}$ l.s.c. proper convex functions, and a matrix $A\in\mathbb{R}^{n\times m}$, let $T_{a}:=T_{a}(A\operatorname{dom}(f),\operatorname{dom}(g))$. Then
@@
The result follows from Theorem 5.1 with the use of [rockafellarConvexAnalysis1970, Corollary 31.2.1]. ∎
@@
Consider 5.1, and let $T_{a}:=T_{a}(A\operatorname{dom}(f),\operatorname{dom}(g))$. A point $x\in\mathbb{R}^{n}$ is a minimizer of 5.1 if, and only if, the following conditions hold:
@@
From Theorem 5.1, the result follows from application of [rockafellarConvexAnalysis1970, Corollary 31.3.1] to the functions $f+\iota_{A^{-1}(T_{a})}$ and $g+\iota_{T_{a}}$. ∎
Infeasible case as convention
In the above results, the domains of the function under consideration may not intersect, in which case $f+g\circ A\equiv\infty$. Intersection is required for the existence of a meaningful joint supporting subspace, so it might be reasonable to make this non-trivial intersection a requirement. This kind of requirement was dubbed a “universal qualification condition” in the context of conic programming [tuncelStrongDualityMinimal2012]. We observe that such a condition is not a true limitation of the theory, because the behaviour outside this condition is known and simple. Therefore, we instead use a convention: whenever $C\cap D=\emptyset$, we let $T_{a}(C,D):=\emptyset$ as well. As a result, $f+\iota_{A^{-1}(T_{a})}\equiv\infty$, and the equality in Corollary 5.5 trivially holds because both sides evaluate to $\infty$, with the value of $\infty$ attained in the dual.
5.3.1 Facial reduction
We derive two known facial reduction results, thereby showing that the joint supporting subspace is indeed a generalization of facial reduction.
Problem 5.2 (Conic program).
For $K\subseteq\mathbb{R}^{n}$ a closed convex cone, $A\in\mathbb{R}^{m\times n}$ a matrix, $b\in\mathbb{R}^{m}$, and $c\in\mathbb{R}^{n}$,
$\mathop{\mathrm{minimize}}_{x\in\mathbb{R}^{n}}\langle x,c\rangle\quad\mathrm{s.t.}\quad Ax=b,\quad x\in K.$
We apply Corollary 5.5 and Corollary 5.6 to this problem, yielding analogues of [tuncelStrongDualityMinimal2012, Theorem 4.3] and [tuncelStrongDualityMinimal2012, Corollary 4.4]. The derived results differ from those in [tuncelStrongDualityMinimal2012] only in that our primal problem is their dual.
Corollary 5.7 (Exact conic dual).
For $A\in\mathbb{R}^{m\times n}$, $b\in\mathbb{R}^{m}$, $c\in\mathbb{R}^{n}$, and a closed convex cone $K\subseteq\mathbb{R}^{n}$,
$\inf\{\langle x,c\rangle\mid x\in\mathbb{R}^{n},Ax=b,\text{ and }x\in K\}$
$\max\{\langle b,y\rangle\mid y\in\mathbb{R}^{m}\text{ and }A^{T}y-c\in F^{\circ}\},$
where $F=K\cap A^{-1}(H_{AK}(\{b\}))$.
Proof of Corollary 5.7.
We use Corollary 5.5 with $f(x):=\langle c,x\rangle+\iota_{K}(x)$, and $g:=\iota_{\{b\}}$. We find
$T:=T(A\operatorname{dom}(f),\operatorname{dom}(g))=T(AK,\{b\})=H_{AK}(\{b\}),$
where the last equality follows from Theorem 4.1. Since $AK$ is a cone, any face includes the origin, and therefore $0\in H_{AK}(\{b\})=T_{a}$, which means that $T_{a}=T$. The functions in the dual program are
$f(x)+\iota_{A^{-1}(T)}(x)$ $=\langle c,x\rangle+\iota_{K}(x)+\iota_{A^{-1}(H_{AK}(\{b\}))}(x)$
$=\langle c,x\rangle+\iota_{K\cap A^{-1}(H_{AK}(\{b\}))}(x)$
and
$g(Ax)+\iota_{T}(Ax)=\iota_{\{b\}}(Ax)+\iota_{H_{AK}(\{b\})}(Ax)=\iota_{\{b\}}(Ax).$
Let $F=K\cap A^{-1}(H_{AK}(\{b\}))$, the subset of $K$ selected by the joint supporting subspace. The set $F$ is a cone, because $H_{AK}(\{b\})\cap AK$ is a face of the cone $AK$ by Definition 6.5 and Corollary 6.1, and $F$ is therefore itself a cone.
Then with [rockafellarConvexAnalysis1970, Theorem 12.3, Theorem 14.1] we compute that
$(f+\iota_{A^{-1}(T)})^{*}(A^{T}y)+(g+\iota_{T})^{*}(-y)$ $=(\langle c,\cdot\rangle+\iota_{F})^{*}(A^{T}y)+(\iota_{\{b\}})^{*}(-y)$
$=\iota_{F^{\circ}}(A^{T}y-c)-\langle b,y\rangle.$
We thereby find the dual problem, for which, by Corollary 5.5, strong duality holds, and the dual is attained. ∎
Corollary 5.8 (Conic KKT conditions).
A point $x\in\mathbb{R}^{n}$ is a minimizer of 5.2 if, and only if, there exists some $y\in\mathbb{R}^{m}$ such that all of the following hold.
  1. 1.
    $Ax=b$.
  2. 2.
    $A^{T}y-c\in F^{\circ}$.
  3. 3.
    $\langle x,A^{T}y-c\rangle=0$.
Proof of Corollary 5.8.
We apply Corollary 5.6 to the case where $f(x):=\langle c,x\rangle+\iota_{K}(x)$ and $g:=\iota_{\{b\}}$. As in the proof of Corollary 5.7, we find that $f(x)+\iota_{A^{-1}(T_{a})}=\langle c,x\rangle+\iota_{F}(x)$ and $g(\bar{y})+\iota_{T_{a}}(\bar{y})=\iota_{\{b\}}(\bar{y})$.
Then $(g+\iota_{T_{a}})^{*}(y)=\langle b,y\rangle$, and the first KKT condition, that $\partial(g+\iota_{T_{a}})^{*}(y)=\{b\}$ reduces to $Ax=b$. The second KKT condition is
$A^{T}y$ $\in\partial(f+\iota_{A^{-1}(T_{a})})(x)$
$=\partial(\langle c,\cdot\rangle+\iota_{F})(x)=c+N_{F}(x).$
But since $N_{F}(x)=F^{\circ}\cap x^{\perp}$, this second condition holds if, and only if, both $x\perp A^{T}y-c$ and $A^{T}y-c\in F^{\circ}$. ∎
@@ 6 A framework for local face lattices
Inspired by the lexicographic characterization of convex faces of [martinez-legaz_LEXICOGRAPHICALCHARACTERIZATIONFACES_], we develop a framework describing the lattice of convex faces at a point $x\in C$, which we will then use to show the results of Section 4. We begin with a characterization of faces of convex sets. In the next few results, we denote $X$ for a real vector space.
Theorem 6.1 (Characterization of convex faces by supporting subspaces).
@@
Theorem 6.1 motivates the next mathematical object.
@@
Though supporting subspaces are affine subspaces, we omit the adjective “affine” for brevity, mirroring the term “supporting hyperplane”. The notion of supporting subspaces seems to be surprisingly sparse in the literature; the only previous appearance of this notion the author could find is in [brondstedIntroductionConvexPolytopes1983, Exercise 5.4]. This idea is of central importance to this paper; it allows us to shift our focus away from the faces themselves to the comparatively simpler affine subspaces which select them through intersections with the convex set. The next result describes the relationship between supporting subspaces and faces.
Corollary 6.1 (Relation of supporting subspaces to faces).
Let $C\subseteq X$ be convex and $S\subseteq C$ non-empty.
@@
Proof of Corollary 6.1.
1) ($\implies$) We show that if $L$ is a supporting subspace, then $F:=C\cap L$ satisfies both conditions of Theorem 6.1:
@@
2) ($\implies$) If $F$ is a face, then we argue that $\operatorname{aff}(F)$ is a suitable supporting subspace. Indeed, by Theorem 6.1 we know that $F=C\cap\operatorname{aff}(F)$ and that $C\setminus F$ is convex. Note that $C\setminus F=C\setminus(C\cap\operatorname{aff}(F))=C\setminus\operatorname{aff}(F)$, and therefore $C\setminus\operatorname{aff}(F)$ is convex, while $F=C\setminus\operatorname{aff}(F)$, which shows that $\operatorname{aff}(F)$ is a supporting subspace that selects $F$ by intersection.
@@
In this work, we focus on “local” supporting subspaces; the subset of supporting subspaces that contain a point $x\in C$. These subsets have some additional nice algebraic properties. We find it useful to generalize this idea of “locality” to supporting subspaces that are localized at a set $S\subseteq C$ instead of only at points, i.e., we consider all supporting subspaces of $C$ that also contain $S$. To emphasize this connection, we say that a supporting subspace is ”at SS” to describe its containment of $S$. The reader should keep in mind the important special case of $S$ being a singleton $S=\{x\}$, for which the results are localized at a point.
@@
When $S$ is the singleton $\{x\}$, the requirement that $S-S\subseteq U$ is trivially satisfied. When $S$ is not a singleton, this assumption guarantees that $\forall s_{1},s_{2}\in S,U+s_{1}=U+s_{2}$. This implies that $\forall s\in S,U+s=U+S$, and so the Minkowski sum $U+S$ is a simple affine offset of $U$, for which $S\subseteq U+S$. In other words, the supporting subspace $U+S$ contains $S$.
Next, we introduce a generalization of the notion of a normal cone. A normal cone of a convex set $C\subseteq\mathbb{R}^{n}$ is defined as
$N_{C}(x):=\{z\in\mathbb{R}^{n}\mid\forall y\in C\langle y-x,z\rangle\leq 0\}.$
When $C$ is a cone and $x=0$, the normal cone $N_{C}(0)=:C^{\circ}$ is called the polar of $C$. The polar is a dual operation, in the sense that
${C^{\circ}}^{\circ}=\operatorname{cl}(C).$
We can see that a normal cone is a dual object which “describes” the original cone only up to its closure. We require a dual object which has sufficient descriptive power to go beyond the closure, by encoding information about the boundary. To this end, we introduce the notion of a nested normal cone. Let $C\subseteq\mathbb{R}^{n}$ be convex and $U\subseteq\mathbb{R}^{n}$ be a subspace, then $\forall x\in\mathbb{R}^{n}$,
@@
The hyperplanes corresponding to the vectors in this normal cone are supporting hyperplanes of $C$ that contain $S$. Note also that $N_{C}(S)=\bigcap_{s\in S}N_{C}(s).$ The idea of normal cones that are located at sets appears in the earlier work [lu_NormalFansPolyhedral_2008], where normal cones are defined for faces of polytopes, instead of at points. Combining the two generalizations, we get the following.
@@
The nested normals at a set are orthogonal with that set.
@@
As previously discussed, one motivating property of the nested normals is that they can encode information about the boundary of convex sets. We make this idea concrete.
The tangent cone of a convex set $C$ at $x\in C$ is the smallest closed cone $T$ such that $C\subseteq x+T$. We can think of this cone as a local description of $C$ around $x$. The normal cone is the polar of the tangent cone, and therefore “describes” the convex set $C$ around $x$ to the same extent as the tangent cone. In contrast, nested normals in supporting subspaces of $C$ at $x$ describe the cone of feasible directions.
Definition 6.4 (Cone of feasible directions).
For a convex set $C\in\mathbb{R}^{n}$ with $x\in C$, we define the cone of feasible directions of $C$ at $x$ to be
$D_{C}(x):=\operatorname{cone}(C-x).$
The closure of the cone of feasible directions is the tangent cone. The difference is that the cone of feasible directions captures additional information about the boundary of $C$ around $x$. The following result shows that any convex cone (including the cone of feasible directions at any point) can be recovered from the nested normals in supporting subspaces.
Theorem 6.2 (Local polar without closure).
Given any convex cone $T$ with $0\in T$,
$T=\{x\in\mathbb{R}^{n}\mid\forall U\in\mathcal{F}(T;0),v\in N_{T}^{U}(0),x\in U\implies\langle x,v\rangle\leq 0\}.$
Proof of Theorem 6.2.
Let $T^{\prime}:=\{x\in\mathbb{R}^{n}\mid\forall U\in\mathcal{F}(T;0),v\in N_{T}^{U}(0),x\in U\implies\langle x,v\rangle\leq 0\}$. By the definition of nested normals, $T\subseteq T^{\prime}$. For the containment ($\supseteq$), we argue by contradiction: we assume that there is some $x\in T^{\prime}\setminus T$. If $x\in\mathrm{int}(T^{c})$, then $x$ is strictly separated from $T$ by a supporting hyperplane of $T$ (e.g., see [rockafellarConvexAnalysis1970, Theorem 11.3]). Any such hyperplane induces a nested normal of the form $v\in N_{T}^{\mathbb{R}^{n}}(0)$, for which $v^{\perp}$ is the hyperplane, and $\langle v,x\rangle>0$. But then $x\notin T^{\prime}$, and so $x\notin\mathrm{int}(T^{c})$. Since we assumed that $x\notin T$, the only possibility is that $x\in\partial T$. Then $x$ is contained in a supporting hyperplane, so there is some nested normal $v\in N_{T}^{\mathbb{R}^{n}}(0)$ such that $\langle v,x\rangle=0$. Let $T_{1}=T\cap v^{\perp}$, and narrow the ambient space to the subspace $T_{1}$. Inside $T_{1}$, we repeat the same argument, and keep narrowing the ambient space in this manner. This process must eventually reduce the ambient space to $\operatorname{\mathrm{span}}(x)$, because $\mathbb{R}^{n}$ is finite-dimensional. In this one-dimensional space, with the narrowed cone $T_{\ell}=T\cap\operatorname{\mathrm{span}}(x)$, note that $T_{\ell}$ is a closed cone. Therefore, either $x\in T_{\ell}$ or $x\in\mathrm{int}(T_{\ell}^{c})$, both of which yields a contradiction, from which the result follows. ∎
Another important property of nested normals is that they describe the containment structure of supporting subspaces, which is closely related to the lattice structure of the faces. We make this precise in the next two results.
@@
The r.h.s. is an exposed face of the face $C\cap(V+S)$, so it is itself a face, which by Corollary 6.1 implies that $V\cap v^{\perp}\in\mathcal{F}(C;S)$. ∎
@@
This result is similar to the lexicographic characterization of the faces of convex sets in [martinez-legaz_LEXICOGRAPHICALCHARACTERIZATIONFACES_, Proposition 6], which states that every face can be obtained by a sequence of faces where each face is an exposed face of the preceding face. Let us state a technical lemma that we will need for the proof.

### v5

@@ 3 Notation and Preliminaries
Let $C\subseteq\mathbb{R}^{n}$ be convex. We denote by $\operatorname{relint}(C)$ the relative interior of $C$. A subset $F\subseteq C$ is defined to be a face of $C$ when $\forall y,z\in C$, $(\exists x\in]y,z[\cap F\implies y,z\in F)$. That is, a face $F\subseteq C$ is a subset of $C$ for which there is no line segment that “goes through” $F$ while remaining in $C$. A face $F\subseteq C$ is exposed when there is a $u\in\mathbb{R}^{n}\setminus\{0\}$ such that $F=\operatorname*{\arg\max}_{c\in C}\langle c,u\rangle$.
@@
For $x\in\mathbb{R}^{n}$ and a convex set $C\subseteq\mathbb{R}^{n}$, the normal cone of $C$ at $x$ is $N_{C}(x)=\{y\mid\forall c\in C,\langle y,c-x\rangle\leq 0\}$ if $x\in C$, and $N_{C}(x)=\emptyset$ otherwise. We denote $\mathbb{R}_{+}:=[0,\infty)$, $\mathbb{R}_{++}:=(0,\infty)$, and similarly for $\mathbb{R}_{-}$ and $\mathbb{R}_{--}$. The extended reals are $\bar{\mathbb{R}}=\mathbb{R}\cup\{-\infty,\infty\}$. Given a convex function $f:\mathbb{R}^{n}\to\bar{\mathbb{R}}$ we denote by $\operatorname{dom}(f)$ the effective domain of $f$; $\operatorname{dom}(f):=\{x\in\mathbb{R}^{n}\mid f(x)\neq\infty\}$. The function $f$ is proper if it is nowhere -$\infty$ and has nonempty effective domain. The subdifferential is
@@
with $\partial f(x)=\emptyset$ when $x\notin\operatorname{dom}(f)$. The convex conjugate is $f^{*}(y):=\sup_{x\in\mathbb{R}^{n}}\langle y,x\rangle-f(x)$.
@@ 4 The joint supporting subspace
We note that when $C\cap D=\emptyset$, we have $T_{a}=\emptyset$ also. This is an intentional convention which takes care of the infeasible case in Section 5. The Minkowski sum $(+C\cap D)$ in the definition of $T_{a}$ is in fact an affine translation. This follows from the next corollary.
@@
It was shown in [23, Theorem 11.3] that this notion is equivalent to a lack of proper separation of the domains. We define the notions of separation, and proper separation below. We define these notions in affine subspaces, an intuitive generalization of separation in Euclidean spaces.
@@
Proper separation excludes the case where the two sets are merely separated by virtue of being both contained in a single hyperplane. Since separation follows from proper separation, and lack of proper separation is equivalent to intersection of relative interiors (by [23, Theorem 11.3]), the contrapositive statement is that
@@
The result follows from Theorem 4.2 by Equation 4, with the caveat that we defined separation in affine subspaces, which are not necessarily vector spaces, while in [23, theorem 11.3] this notion is defined in $\mathbb{R}^{n}$. Yet the argument still holds because the notions of separation, proper separation, and relative interiors are invariant under translation of $T_{a}$ to $T$. For any point $x_{0}\in T_{a}$, $C\cap T_{a}$ is (properly) separated from $C\cap T_{a}$ iff $C\cap T_{a}-x_{0}$ is (properly) separated from $D\cap T_{a}-x_{0}$ in $T$. A similar equivalence holds for intersection of relative interiors. ∎
@@ 5 Main results
With Corollary 4.5 we provide qualification-free generalizations by localizing to $T_{a}$. The next result shows that when standard qualification conditions hold, the classical formulations are trivially recovered.
@@ 5.1 Subdifferential calculus
We note that in the infeasible case, when $f+g\equiv\infty$, $T_{a}$ is $\emptyset$, and the result holds trivially. A similar pattern occurs for many results in this section.
@@
By Corollary 4.5 applied to $\operatorname{dom}(f)$ and $\operatorname{dom}(g)$, the result follows by application of [23, Theorem 23.8] to the functions $f+\iota_{T_{a}}$ and $g+\iota_{T_{a}}$. ∎
@@
With Corollary 4.5 we can use [23, Corollary 23.8.1], from which we find that
@@
If $\operatorname{dom}(g)\cap\operatorname{\mathrm{range}}(A)=\emptyset,$ then the result holds trivially because $g(Ax)\equiv\infty$, and $(g+\iota_{T_{a}})(Ax)\equiv\infty$ as well. Otherwise, by Corollary 4.5, $\operatorname{relint}\operatorname{dom}(g+\iota_{T_{a}})\cap\operatorname{\mathrm{range}}(A)\neq\emptyset$, which lets us apply [23, Theorem 23.9] to the composition $g+\iota_{T_{a}}]\circ A$. Additionally, note that the only face of $\operatorname{\mathrm{range}}(A)$ is itself, whence by Lemma 4.1 $\operatorname{\mathrm{range}}(A)\subseteq T_{a}$. Since $0\in\operatorname{\mathrm{range}}(A)$, this in fact implies that $T_{a}=T$, and the result follows. ∎
@@ 5.2 Infimal convolution
Whenever $\operatorname{dom}(f)\cap\operatorname{dom}(g)=\emptyset$, we have that $T_{a}=\emptyset$, in which case the equality holds trivially, with the infimum attained at any $y\in\mathbb{R}^{n}$. Otherwise, by Corollary 4.5 the result follows from [23, Theorem 20.1]. ∎
@@ 5.3 Convex optimization
which is a weakening of Slater’s condition. A program with this property is called strongly consistent [23, Theorem 31.2].
@@
For $f:\mathbb{R}^{n}\to\bar{\mathbb{R}}$ and $g:\mathbb{R}^{m}\to\bar{\mathbb{R}}$ lower semicontinuous proper convex functions, a matrix $A\in\mathbb{R}^{m\times n}$, $T_{a}:=T_{a}(A\operatorname{dom}(f),\operatorname{dom}(g))$, and any $x\in\mathbb{R}^{n}$,
@@
where Equation 7 holds by [23, Theorem 6.6]; Equation 9 uses the identity that $A(S_{1}\cap A^{-1}(S_{2}))=AS_{1}\cap S_{2}$ for any sets $S_{1}\subseteq\mathbb{R}^{n}$, $S_{2}\subseteq\mathbb{R}^{m}$; and that the set in Equation 9 is non-empty is guaranteed by Corollary 4.5. ∎
@@
For $f:\mathbb{R}^{n}\to\bar{\mathbb{R}}$ and $g:\mathbb{R}^{m}\to\bar{\mathbb{R}}$ l.s.c. proper convex functions, and a matrix $A\in\mathbb{R}^{m\times n}$, let $T_{a}:=T_{a}(A\operatorname{dom}(f),\operatorname{dom}(g))$. Then
@@
The result follows from Theorem 5.1 with the use of [23, Corollary 31.2.1]. ∎
@@
Consider 5.1, and let $T_{a}:=T_{a}(A\operatorname{dom}(f),\operatorname{dom}(g))$. A point $x\in\mathbb{R}^{n}$ is a minimizer of 5.1 if, and only if, there exists some $y\in\mathbb{R}^{m}$ such that
@@
From Theorem 5.1, the result follows from application of [23, Corollary 31.3.1] to the functions $f+\iota_{A^{-1}(T_{a})}$ and $g+\iota_{T_{a}}$. ∎
@@ 6 A framework for local face lattices
Inspired by the lexicographic characterization of convex faces of [16], we develop a framework describing the lattice of convex faces at a point $x\in C$, which we will then use to show the results of Section 4. We begin with a characterization of faces of convex sets. In the next few results, we denote $X$ for a real vector space.
Proposition 6.1 (Characterization of convex faces by supporting subspaces).
@@
Proposition 6.1 motivates the next mathematical object.
@@
Though supporting subspaces are affine subspaces, we omit the adjective “affine” for brevity. This notion appears in [6, Exercise 5.4], and is related to order ideals, a concept defined for convex cones. Supporting subspaces are of central importance to this paper; it allows us to shift our focus away from the faces themselves to the comparatively simpler affine subspaces which select them through intersections with the convex set. The next result describes the relationship between supporting subspaces and faces.
Proposition 6.2 (Relation of supporting subspaces to faces).
Let $C\subseteq X$ be convex.
@@
Proof of Proposition 6.2.
1) ($\implies$) We show that if $L$ is a supporting subspace, then $F:=C\cap L$ satisfies both conditions of Proposition 6.1:
@@
2) ($\implies$) If $F$ is a face, then we argue that $\operatorname{aff}(F)$ is a suitable supporting subspace. Indeed, by Proposition 6.1 we know that $F=C\cap\operatorname{aff}(F)$ and that $C\setminus F$ is convex. Note that $C\setminus F=C\setminus(C\cap\operatorname{aff}(F))=C\setminus\operatorname{aff}(F)$, and therefore $C\setminus\operatorname{aff}(F)$ is convex, while $F=C\setminus\operatorname{aff}(F)$, which shows that $\operatorname{aff}(F)$ is a supporting subspace that selects $F$ by intersection.
@@
In this work, we focus on “local” supporting subspaces; the subset of supporting subspaces that contain a point $x\in C$. These subsets have some additional nice algebraic properties. We find it useful to generalize this idea of “locality” to supporting subspaces that are localized at a set $S\subseteq C$ instead of only at points, i.e., we consider all supporting subspaces of $C$ that also contain $S$. To emphasize this connection, we say that a supporting subspace is “at SS” to describe its containment of $S$. The reader should keep in mind the important special case of $S$ being a singleton $S=\{x\}$, for which the results are localized at a point.
@@
When $S$ is the singleton $\{x\}$, the requirement that $S-S\subseteq U$ is trivially satisfied. When $S$ is not a singleton, this assumption guarantees that $\forall s_{1},s_{2}\in S,U+s_{1}=U+s_{2}$. This implies that $\forall s\in S,U+s=U+S$, and so the Minkowski sum $U+S$ is a simple affine offset of $U$, for which $S\subseteq U+S$. In other words, the supporting subspace $U+S$ contains $S$. We introduce the notion of a nested normal cone. Let $C\subseteq\mathbb{R}^{n}$ be convex and $U\subseteq\mathbb{R}^{n}$ be a subspace, then $\forall x\in\mathbb{R}^{n}$,
@@
The hyperplanes corresponding to the vectors in this normal cone are supporting hyperplanes of $C$ that contain $S$. Note also that $N_{C}(S)=\bigcap_{s\in S}N_{C}(s).$ The idea of normal cones that are located at sets appears in the earlier work [13], where normal cones are defined for faces of polytopes, instead of at points. Combining the two generalizations, we get the following.
@@
Next, we show that the nested normals at a set $S$ are orthogonal with SSS-S.
@@
The core property of nested normals is that they describe the lattice structure of supporting subspaces, which is closely related to the lattice structure of the faces. We make this idea precise in the next two results.
@@
The r.h.s. is an exposed face of the face $C\cap(V+S)$, so it is itself a face, which by Proposition 6.2 implies that $V\cap v^{\perp}\in\mathcal{F}(C;S)$. ∎
@@
This result is similar to the lexicographic characterization of the faces of convex sets in [16, Proposition 6], which states that every face can be obtained by a sequence of faces where each face is an exposed face of the preceding face. Let us state a technical lemma that we will need for the proof.